# N Hale, E Thomann, JAC Weideman

### v1

@@ Abstract
Functional differential equation, bifurcation, continuation, Laguerre spectral collocation
@@ 1 Introduction
$u^{\prime}(t)+a\,u(t)=b\,u(\alpha t),\quad t>0,$ (1.1)
with $a$, $b$, and $\alpha>0$ constants, was the topic of two major papers, (Fox et al., 1971) and (Kato and McLeod, 1971), both dating back about half a century. An application mentioned in the former of these references is the dynamics of an overhead current collection system for an electric locomotive. Even earlier papers derived the equation as a probabilistic model of absorption processes (Ambartsumian, 1944; Gaver, 1964). In later years the equation has re-appeared in mathematical biology, such as in the modeling of cell growth (Hall and Wake, 1989; van Brunt and Vlieg-Hulstman, 2010a).
@@
$u^{\prime}(t)+u(t)=u^{2}(\alpha t),\quad t>0,$ (1.2)
@@
considered by (Athreya, 1985) and more recently by (Dascaliuc et al., 2019). The latter paper, in which the term $\alpha$-Riccati equation is coined, discusses an application to cascading processes in Fourier space representations of some nonlinear PDEs (including the Navier–Stokes equations). The two references cited both approach the nonlinear problem from a probabilistic viewpoint, in particular a stochastic branching process. Here, the tool is numerical computation supported by a perturbation analysis.
@@
For $\alpha=1$, both (1.1) and (1.2) reduce to well-known equations that will not be considered here. For $\alpha<1$, the equations are delay differential equations with proportional or ‘pantograph-type’ delay. This regime was investigated in (Fox et al., 1971) and (Kato and McLeod, 1971) in the linear case, and (Dascaliuc et al., 2019) in the nonlinear case. In particular, existence and uniqueness of solutions were established for both (1.1) and (1.2).
@@
The case $\alpha>1$ was considered by (Fox et al., 1971) and (Kato and McLeod, 1971) for equation (1.1), but for the nonlinear problem (1.2) many gaps in current knowledge remain. A status report can be seen in (Dascaliuc et al., 2019, Fig. 2), a diagram that catalogs the existence/uniqueness properties in several regions of the $(\alpha,u_{0})$ plane, with $u(0)=u_{0}$ the initial condition. In several sizeable regions either existence or uniqueness is marked as “unknown.” The focus here will be on solutions to (1.2) for the case $\alpha>1$, $u_{0}=1$, which forms a small but rather important section of that diagram.
@@
We shall make no attempt to prove any theorems regarding existence/uniqueness, but rather focus on quantitative aspects such as the approximation of solutions, or, since multiple solutions exist, to determine how many solutions in a certain class are possible for each value of $\alpha$. By a combination of analysis and computation we shall identify parameter regions where the solutions to the nonlinear problem (1.2) can be approximated well by certain explicitly known series solutions of the linear problem (1.1).
@@
For numerical solutions we use a spectral collocation method for functional differential equations. (Details can be found in (Hale, 2023a) and a short description in section 4.) This method provides more efficient and more accurate results than other methods reported in the literature. Two such methods for the linear problem (1.1) are described in (Fox et al., 1971): A finite difference method, applicable to the case $\alpha<1$, and a Lanczos tau-method for both $\alpha<1$ and $\alpha>1$. Numerical results, of low accuracy in today’s terms but impressive for 1970s technology, are given. The only numerical approach to the nonlinear problem (1.2) that we are aware of is the probabilistic simulation of (Dascaliuc et al., 2023). Like most methods of Monte Carlo-type, however, it is not efficient for doing the extensive experimentation that will be reported here. Because the problem (1.2) is nonlinear, any numerical procedure would require iteration and this demands good starting guesses. Such approximations, which may be of interest in their own right, are obtained here by an elementary perturbation analysis.
@@
The outline of this paper is therefore as follows. In section 2 we introduce some aspects of the linear problem (1.1) that are pertinent in the analysis of the nonlinear problem (1.2). In particular, we consider $a=1$ and $b=2$, corresponding to the linearisation of (1.2) about the constant solution $u=1$, and demonstrate the existence of nontrivial solutions to the homogeneous linear problem for certain characteristic values of $\alpha$, designated here as $\alpha_{n}$, $n=1,2,\ldots$. In section 3 we show that these characteristic solutions, appropriately scaled, provide the leading order term in a perturbation of the constant solution to (1.2) in the neighbourhood of these characteristic values. In section 4 we use these perturbation approximations in combination with the spectral collocation method and pseudo-arclength continuation to determine numerically a family of solutions to (1.2) for a range of $\alpha$ values, which we discuss in section 5.
@@ 2 Summary of results for the linear equation
The papers (Fox et al., 1971) and (Kato and McLeod, 1971) contain a wealth of information on the linear equation (1.1). This section highlights only the major results that will be used in the nonlinear analysis of the next section. In particular, in that section the values $a=1$ and $b=2$ will be relevant, i.e.,
$u^{\prime}(t)+u(t)=2u(\alpha t).$ (2.1)
@@
$u(t)=CE(t;\alpha),\quad\text{where}\quad E(t;\alpha)=e^{-t}\Big{(}1+\sum_{k=1}^{\infty}\frac{2^{k}e^{(1-\alpha^{k})t}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}\Big{)},$ (2.2)
and $C$ is a constant. This solution is a special case of one given in (Kato and McLeod, 1971) for (1.1). [^1] Neither (Kato and McLeod, 1971) nor (Fox et al., 1971) explicitly derive (2.2) but (Hall and Wake, 1989) and (van Brunt and Vlieg-Hulstman, 2010a) contain derivations based on Laplace transforms and Mellin transforms, respectively. The series converges absolutely and uniformly for all $t\geq 0$ in the case $\alpha>1$.
The initial condition $u(0)=u_{0}$ can be matched to (2.2) by setting $t=0$, but of particular interest is the case $u_{0}=0$. This can be satisfied by setting $C=0$, but nontrivial solutions are possible for special values of $\alpha$, as the following lemma shows. These values of $\alpha$ can therefore be viewed as eigenvalues or characteristic values of the linear operator. In other words, for these values of $\alpha$ the function (2.2) lies in the nullspace of (2.1).
@@
$01+\sum_{k=1}^{\infty}\frac{2^{k}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}=$ 0 (2.3)
@@
Proof: An identity well-known in the field of $q$-series (Gasper and Rahman, 2004, p. 11) states that
$11+\sum_{k=1}^{\infty}\frac{(-1)^{k}q^{k(k+1)/2}c^{k}}{(1-q)(1-q^{2})\cdots(1-q^{k})}=(1-cq)(1-cq^{2})\cdots(1-cq^{n})\cdots,\quad|q|<$ 1 . (2.4)
The proof is completed by substituting $c=2$, $q=1/\alpha$. ∎
A related observation was made in (Kato and McLeod, 1971), namely that the initial condition $u(0)=0$ “can occur only for exceptional pairs $a$ and $b$” (with these constants as defined in (1.1)). The authors did not explore further, but in (van Brunt and Vlieg-Hulstman, 2010a, Eq. (2.13)) it was shown that $u(0)=0$ if and only if $b=a\alpha^{n}$ for some integer $n$. This follows from (2.4) by substituting $c=b/a$ and $q=1/\alpha$. Note that both (Kato and McLeod, 1971) and (van Brunt and Vlieg-Hulstman, 2010a) addressed the situation with $\alpha$ fixed; our interest is in the case when $a$ and $b$ are fixed.
To distinguish these cases we shall call the functions (2.2) with $\alpha=\alpha_{n}=2^{1/n}$ “characteristic functions” rather than “eigenfunctions” as in (van Brunt and Vlieg-Hulstman, 2010a). The first few characteristic functions of (2.1) are shown in Fig. 1, normalised such that $\|u\|_{2}=1$ (the norm is the standard, unweighted $L^{2}$ norm on $[0,\infty)$.)
Figure 1: $L^{2}$ normalised characteristic solutions to (2.1), as defined by the series (2.2), for $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots,6$. An additional scaling of $(-1)^{n+1}$ is also included so that all solutions are initially positive.
Noteworthy features of the characteristic solutions shown in Fig. 1 include the observation that for $n=1$ the function has no zero in $(0,\infty)$, for $n=2$ there is one zero, for $n=3$ two zeros, and so on. This is reminiscent of Sturm–Liouville theory, but here the roots of two successive characteristic functions do not strictly interlace. (See also (van Brunt and Vlieg-Hulstman, 2010b) for comments about orthogonality, but keep in mind that their definition of an eigenfunction is different, as mentioned below Lemma 1.) Another striking feature is the very flat profile near $t=0$. By repeated differentiation of (2.1) one can show that if $u(0)=0$, then all derivatives of $u$ vanish at $t=0$. This is suggestive of an essential singularity at the origin, which is consistent with the asymptotic estimate of $u$ as $t\to 0^{+}$ given in (Fox et al., 1971, Thm. 3). The asymptotic behaviour as $t\to+\infty$, on the other hand, is given in (Fox et al., 1971, Thm. 2), namely, $u={\rm O}(e^{-t})$ as $t\to+\infty$. (This follows from (2.1) under the assumption that $u(\alpha t)\ll u(t)$ for large $t$.)
@@ 3 Perturbation analysis of the nonlinear equation
As mentioned in the introduction, the focus of this paper is on the nonlinear problem (1.2) with initial condition $u(0)=1$. In (Dascaliuc et al., 2019, Thm. 4.1) two families of global solutions have been identified according to the asymptotic limits (A) $u\to 1$ as $t\to\infty$ and (B) $u\to 0$ as $t\to\infty$. Here we restrict ourselves to the cases where these two limits are approached exponentially, i.e., according to (A) $u=1+{\rm O}(e^{-t})$ and (B) $u={\rm O}(e^{-t})$. Algebraic approaches to these limits may also be possible (Dascaliuc et al., 2024), but this will require a different procedure from the one described here.
@@
Because our focus is on solutions in class (A), define $v(t)=u(t)-1$. Substituting into (1.2) gives
$v^{\prime}(t)+v(t)=2v(\alpha t)+v^{2}(\alpha t),\quad v(0)=0.$ (3.1)
With the intuition (and support from numerical experimentation) that the characteristic values from the previous section act as bifurcation points in the nonlinear equations (1.2) and (3.1), we seek solutions in the neighbourhood of these points by considering $\alpha=\alpha_{n}+\epsilon$, $|\epsilon|\ll 1$. Further, assuming small perturbations about the trivial solution $v(t)=0$, we try the regular expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$. This gives, to leading order,
$v_{0}^{\prime}(t)+v_{0}(t)=2v_{0}(\alpha_{n}t),\quad v_{0}(0)=0,$ (3.2)
which is the linear equation discussed in section 2.222In an effort to avoid overencumbering notation, we omit burdening $v_{0}$ with an additional subscript, $n$. However, it should be understood here, and throughout, that each $v_{0}$ corresponds to a particular $\alpha=\alpha_{n}$. This value of $n$ should be clear from context. By Lemma 1, the explicit solution to this equation is given by $v_{0}=CE(t;\alpha_{n})$, where $E(t;\alpha_{n})$ is defined in (2.2). The constant $C$ remains to be determined, however, and this can be done by matching solutions of the linear and nonlinear problems.
To do this, we consider the moments of $v$ as follows. For arbitrary $\alpha$, integrate (3.1) with respect to $t$ to obtain
@@
Consider now the particular perturbation solution corresponding to $n=1$, so that $\alpha=2+\epsilon$. Inserting this and $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$ into (3.4) and then collecting terms gives, at leading order,
@@
By plugging in the series solution $v_{0}=C_{1}E(t;2)$ from (2.2) one finds that the required value of $C_{1}$ is
@@
where the numerical value has been computed by the formulas (A.19) in the Appendix. Similar formulas for $C_{n}$, $n>1$, can be derived by considering higher moments of $v$. Their numerical values are given in Table 1 but details of the derivation are deferred to the Appendix.
$n$
$\hskip 16.0ptC_{n}$
$\hskip 13.0pt\|C_{n}E_{n}\|_{2}$
@@
Table 1: The first six scaling coefficients, $C_{1},\ldots,C_{6}$, as given by (A.16), accurate to all digits shown. Although the $C_{n}$ grow rapidly with $n$, the characteristic functions, $E_{n}(t)=E(t,\alpha_{n})$, are at the same time decreasing in magnitude. Their product, which is the coefficient of $\epsilon$ in the perturbation approximation (3.7), also increases in magnitude, but not as rapidly as $C_{n}$. Nevertheless, this means that as $\alpha$ approaches the value $1$ with $\epsilon$ fixed, the perturbation approximations become less accurate.
In summary, the approximation to (3.1) corresponding to $\alpha=2^{1/n}+\epsilon$ is given to first order by
$v(t)\approx C_{n}E(t;2^{1/n})\epsilon,$ (3.7)
where $E(t;2^{1/n})$ is the series solution (2.2) and the $C_{n}$ values are given in Table 1.
We have not pursued a formal proof of the validity of (3.7) in any asymptotic sense. (The analysis is rather complicated and perhaps of secondary importance because the principal use of (3.7) is to provide initial guesses for the numerical method.) Instead, here is a computational check. We substitute (3.7) into (3.1), and, expecting the remainder to be ${\rm O}(\epsilon^{2})$, we write
$v^{\prime}(t)+v(t)-2v((2^{1/n}+\epsilon)t)-v^{2}((2^{1/n}+\epsilon)t)=r_{n}(t;\epsilon)\,\epsilon^{2},$ (3.8)
for some coefficient function, $r_{n}$. Every quantity on the left can be explicitly computed from the series (2.2) and the tabulated values of $C_{n}$. Therefore the values of $r_{n}(t;\epsilon)$ can be computed for each $\epsilon$. By Taylor analysis, the limiting expressions as $\epsilon\to 0$ are found to be
$r_{n}(t;0)=-2C_{n}E^{\prime}(2^{1/n}t;2^{1/n})t-\big{(}C_{n}E(2^{1/n}t;2^{1/n})\big{)}^{2}.$ (3.9)
These functions are shown for a few values of $n$ in Figure 2. Evidently, the $r_{n}(t;0)$ are bounded for all $t\geq 0$, which provides evidence that the approximation (3.7) satisfies (3.1) uniformly to ${\rm O}(\epsilon^{2})$. The implied constant grows as $\alpha\to 1$, however, as can be seen by examining the vertical scales in Figure 2. Therefore, the usefulness of (3.7) as supplier of initial guesses to the numerical method again diminishes accordingly as $n$ increases.
Figure 2: The coefficient functions $r_{n}(t;\epsilon)$ as defined by (3.8), in the limit as $\epsilon\to 0$. The observation that these functions are bounded on $[0,\infty)$ provides support for the validity of the formulas (3.7) as uniform approximations to the solutions of (3.1).
The details of the numerical method are postponed to the next section, but let us now solve (1.2) (via (3.1)) for values of $\alpha$ close to the characteristic values. A few such solutions are shown in Figure 3, where the perturbation and numerical approximations are displayed. Evidently, the perturbation approximations provide good qualitative descriptions of the solutions.
@@
Figure 3: Numerical solutions $v$ of the perturbation equation (3.1) (solid curves) in comparison to the perturbation approximations (3.7) (dashed curves). The scaling constants $C_{1}$, $C_{2}$ and $C_{3}$ are given in Table 1. In each frame the value of $\alpha$ is $\alpha_{n}+\epsilon$, with numerical values of these parameters given in the titles and legends.

### v2

# Approximate solutions to a nonlinear functional differential equation

@@ Abstract
Functional differential equation, bifurcation, Laguerre spectral collocation
{MSCcodes}
34K18, 34K28, 74S25
@@ 1 Introduction
(1) $u^{\prime}(t)+a\,u(t)=b\,u(\alpha t),\quad t>0,$
with $a$, $b$, and $\alpha>0$ constants, was the topic of two major papers, [10] and [16], both dating back about half a century. An application mentioned in the former of these references is the dynamics of an overhead current collection system for an electric locomotive. Even earlier papers derived the equation as a probabilistic model of absorption processes [1, 12]. In later years the equation has re-appeared in mathematical biology, such as in the modeling of cell growth [15, 19].
@@
(2) $u^{\prime}(t)+u(t)=u^{2}(\alpha t),\quad t>0,$
@@
considered by [2] and more recently by [7]. The latter paper, in which the term $\alpha$-Riccati equation is coined, discusses an application to cascading processes in Fourier space representations of some nonlinear PDEs (including the Navier–Stokes equations). The two references cited both approach the nonlinear problem from a probabilistic viewpoint, in particular a stochastic branching process. Here, the tool is numerical computation supported by a perturbation analysis.
@@
For $\alpha=1$, both (1) and (2) reduce to well-known equations that will not be considered here. For $\alpha<1$, the equations are delay differential equations with proportional or ‘pantograph-type’ delay. This regime was investigated in [10] and [16] in the linear case, and [7] in the nonlinear case. In particular, existence and uniqueness of solutions were established for both (1) and (2).
The case of $\alpha>1$ in (1) and (2) involves a ‘time-advanced’ argument, however, in applications this typically represents a non-temporal variable, thus avoiding a violation of causality in the usual time-evolution sense. For example, the independent variable may represent a scale in the biological settings treated in [15], the magnitude of the wave vector in the self-similar Navier–Stokes equations as discussed in [7], or a transparency scale as in [1]. In the pantograph-type systems introduced in [10], and carefully analyzed in [16], the parameter $\alpha$ is also allowed to take values greater than unity for modeling purposes. Further examples in which $\alpha>1$ arise are models of nerve axons from conduction theory [4] and preview control from control theory, where it models anticipation [3]. As the simplest nontrivial nonlinear equation with this feature, we investigate solutions to (2) in order to provide insight into more complex time-advanced functional differential equations arising in applications.
@@
While [10] and [16] give a thorough analysis of properties of solutions of the linear equation (1), current knowledge for the non-linear counterpart (2) is incomplete. A visual status report is given in [7, Fig. 2]. This diagram catalogs the existence and uniqueness properties of solutions in several regions of the $(\alpha,u_{0})$ plane, where $u(0)=u_{0}$ represents an initial condition. In several regions either existence or uniqueness remains unknown. The focus in this manuscript will be on solutions to (2) for the case $\alpha>1,$ $u_{0}=1$, which forms a small but important section of that diagram.
@@
We shall make no attempt to prove any theorems regarding existence/uniqueness, but rather focus on quantitative aspects such as the approximation of solutions, or, since multiple solutions exist, to determine how many solutions in a certain class are possible for each value of $\alpha$. By a combination of analysis and computation we shall identify parameter regions where the solutions to the nonlinear problem (2) can be approximated well by certain explicitly known series solutions of the linear problem (1).
@@
For numerical solutions we use a spectral collocation method for functional differential equations. (Details can be found in [14] and a short description in section 4.) This method provides more efficient and more accurate results than other methods reported in the literature. Two such methods for the linear problem (1) are described in [10]: A finite difference method, applicable to the case $\alpha<1$, and a Lanczos tau-method for both $\alpha<1$ and $\alpha>1$. Numerical results, of low accuracy in today’s terms but impressive for 1970s technology, are given. The only numerical approach to the nonlinear problem (2) that we are aware of is the probabilistic simulation of [5]. Like most methods of Monte Carlo-type, however, it is not efficient for doing the extensive experimentation that will be reported here. Because the problem (2) is nonlinear, any numerical procedure would require iteration and this demands good starting guesses. Such approximations, which may be of interest in their own right, are obtained here by an elementary perturbation analysis.
@@
The outline of this paper is therefore as follows. In section 2 we introduce some aspects of the linear problem (1) that are pertinent in the analysis of the nonlinear problem (2). In particular, we consider $a=1$ and $b=2$, corresponding to the linearisation of (2) about the constant solution $u=1$, and demonstrate the existence of nontrivial solutions to the homogeneous linear problem for certain characteristic values of $\alpha$, designated here as $\alpha_{n}$, $n=1,2,\ldots$. In section 3 we show that these characteristic solutions, appropriately scaled, provide the leading order term in a perturbation of the constant solution to (2) in the neighbourhood of these characteristic values. In section 4 we use these perturbation approximations in combination with the spectral collocation method and pseudo-arclength continuation to determine numerically a family of solutions to (2) for a range of $\alpha$ values, which we discuss in section 5.
@@ 2 Summary of results for the linear equation
The papers [10] and [16] contain a wealth of information on the linear equation (1). This section highlights only the major results that will be used in the nonlinear analysis of the next section. In particular, in that section the values $a=1$ and $b=2$ will be relevant, i.e.,
(3) $u^{\prime}(t)+u(t)=2u(\alpha t).$
@@
(4) $u(t)=CE(t;\alpha),\quad\text{where}\quad E(t;\alpha)=e^{-t}\Big{(}1+\sum_{k=1}^{\infty}\frac{2^{k}e^{(1-\alpha^{k})t}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}\Big{)},$
and $C$ is a constant. This solution is a special case of one given in [16] for (1). [^1] Neither [16] nor [10] explicitly derive (4) but [15] and [19] contain derivations based on Laplace transforms and Mellin transforms, respectively. The series converges absolutely and uniformly for all $t\geq 0$ in the case $\alpha>1$.
The initial condition $u(0)=u_{0}$ can be matched to (4) by setting $t=0$, but of particular interest is the case $u_{0}=0$. This can be satisfied by setting $C=0$, but nontrivial solutions are possible for special values of $\alpha$, as the following lemma shows. These values of $\alpha$ can therefore be viewed as eigenvalues or characteristic values of the linear operator. In other words, for these values of $\alpha$ the function (4) lies in the nullspace of (3).
@@
(5) $01+\sum_{k=1}^{\infty}\frac{2^{k}}{(1-\alpha)(1-\alpha^{2})\cdots(1-\alpha^{k})}=$ 0
@@
Proof: An identity well-known in the field of $q$-series [11, p. 11] states that
(6) $11+\sum_{k=1}^{\infty}\frac{(-1)^{k}q^{k(k+1)/2}c^{k}}{(1-q)(1-q^{2})\cdots(1-q^{k})}=(1-cq)(1-cq^{2})\cdots(1-cq^{n})\cdots,\quad|q|<$ 1 .
The proof is completed by substituting $c=2$, $q=1/\alpha$.
A related observation was made in [16], namely that the initial condition $u(0)=0$ “can occur only for exceptional pairs $a$ and $b$” (with these constants as defined in (1)). The authors did not explore further, but in [19, Eq. (2.13)] it was shown that $u(0)=0$ if and only if $b=a\alpha^{n}$ for some integer $n$. This follows from (6) by substituting $c=b/a$ and $q=1/\alpha$. Note that both [16] and [19] addressed the situation with $\alpha$ fixed; our interest is in the case when $a$ and $b$ are fixed.
To distinguish these cases we shall call the functions (4) with $\alpha=\alpha_{n}=2^{1/n}$ “characteristic functions” rather than “eigenfunctions” as in [19]. The first few characteristic functions of (3) are shown in Fig. 1, normalised such that $\|u\|_{2}=1$ (the norm is the standard, unweighted $L^{2}$ norm on $[0,\infty)$.)
Figure 1: $L^{2}$ normalised characteristic solutions to (3), as defined by the series (4), for $\alpha_{n}=2^{1/n}$, $n=1,2,\ldots,6$. An additional scaling of $(-1)^{n+1}$ is also included so that all solutions are initially positive.
Noteworthy features of the characteristic solutions shown in Fig. 1 include the observation that for $n=1$ the function has no zero in $(0,\infty)$, for $n=2$ there is one zero, for $n=3$ two zeros, and so on. This is reminiscent of Sturm–Liouville theory, but here the roots of two successive characteristic functions do not strictly interlace. (See also [18] for comments about orthogonality, but keep in mind that their definition of an eigenfunction is different, as mentioned below Lemma 1.) Another striking feature is the very flat profile near $t=0$. By repeated differentiation of (3) one can show that if $u(0)=0$, then all derivatives of $u$ vanish at $t=0$. This is suggestive of an essential singularity at the origin, which is consistent with the asymptotic estimate of $u$ as $t\to 0^{+}$ given in [10, Thm. 3]. The asymptotic behaviour as $t\to+\infty$, on the other hand, is given in [10, Thm. 2], namely, $u={\rm O}(e^{-t})$ as $t\to+\infty$. (This follows from (3) under the assumption that $u(\alpha t)\ll u(t)$ for large $t$.)
@@ 3 Perturbation analysis of the nonlinear equation
As mentioned in the introduction, the focus of this paper is on the nonlinear problem (2) with initial condition $u(0)=1$. In [7, Thm. 4.1] two families of global solutions have been identified according to the asymptotic limits (A) $u\to 1$ as $t\to\infty$ and (B) $u\to 0$ as $t\to\infty$. Here we restrict ourselves to the cases where these two limits are approached exponentially, i.e., according to (A) $u=1+{\rm O}(e^{-t})$ and (B) $u={\rm O}(e^{-t})$. Algebraic approaches to these limits may also be possible [6], but this will require a different procedure from the one described here.
@@
Because our focus is on solutions in class (A), we define $v(t)=u(t)-1$. Substituting into (2) gives
(7) $v^{\prime}(t)+v(t)=2v(\alpha t)+v^{2}(\alpha t),\quad v(0)=0.$
With the intuition (and support from numerical experimentation) that the characteristic values from the previous section act as bifurcation points in the nonlinear equations (2) and (7), we seek solutions in the neighbourhood of these points by considering $\alpha=\alpha_{n}+\epsilon$, $|\epsilon|\ll 1$. Further, assuming small perturbations about the trivial solution $v(t)=0$, we try the regular expansion $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$. This gives, to leading order,
(8) $v_{0}^{\prime}(t)+v_{0}(t)=2v_{0}(\alpha_{n}t),\quad v_{0}(0)=0,$
which is the linear equation discussed in section 2.222In an effort to avoid overencumbering notation, we omit burdening $v_{0}$ with an additional subscript, $n$. However, it should be understood here, and throughout, that each $v_{0}$ corresponds to a particular $\alpha=\alpha_{n}$. This value of $n$ should be clear from context. By Lemma 1, the explicit solution to this equation is given by $v_{0}=CE(t;\alpha_{n})$, where $E(t;\alpha_{n})$ is defined in (4). The constant $C$ remains to be determined, however, and this can be done by matching solutions of the linear and nonlinear problems.
To do this, we consider the moments of $v$ as follows. For arbitrary $\alpha$, integrate (7) with respect to $t$ to obtain
@@
Consider now the particular perturbation solution corresponding to $n=1$, so that $\alpha=2+\epsilon$. Inserting this and $v(t)=\epsilon v_{0}(t)+\epsilon^{2}v_{1}(t)+\cdots$ into (10) and then collecting terms gives, at leading order,
@@
By plugging in the series solution $v_{0}=C_{1}E(t;2)$ from (4) one finds that the required value of $C_{1}$ is
@@
where the numerical value has been computed by the formulas (36) in the Appendix. Similar formulas for $C_{n}$, $n>1$, can be derived by considering higher moments of $v$. Their numerical values are given in Table 1 but details of the derivation are deferred to the Appendix.
$n$ $\hskip 16.0ptC_{n}$ $\hskip 13.0pt\|C_{n}E_{n}\|_{2}$
@@
Table 1: The first six scaling coefficients, $C_{1},\ldots,C_{6}$, as given by (33), accurate to all digits shown. Although the $C_{n}$ grow rapidly with $n$, the characteristic functions, $E_{n}(t)=E(t,\alpha_{n})$, are at the same time decreasing in magnitude. Their product, which is the coefficient of $\epsilon$ in the perturbation approximation (13), also increases in magnitude, but not as rapidly as $C_{n}$. Nevertheless, this means that as $\alpha$ approaches the value $1$ with $\epsilon$ fixed, the perturbation approximations become less accurate.
In summary, the approximation to (7) corresponding to $\alpha=2^{1/n}+\epsilon$ is given to first order by
(13) $v(t)\approx C_{n}E(t;2^{1/n})\epsilon,$
where $E(t;2^{1/n})$ is the series solution (4) and the $C_{n}$ values are given in Table 1.
We have not pursued a formal proof of the validity of (13) in any asymptotic sense. (The analysis is rather complicated and perhaps of secondary importance because the principal use of (13) is to provide initial guesses for the numerical method.) Instead, here is a computational check. We substitute (13) into (7), and, expecting the remainder to be ${\rm O}(\epsilon^{2})$, we write
(14) $v^{\prime}(t)+v(t)-2v((2^{1/n}+\epsilon)t)-v^{2}((2^{1/n}+\epsilon)t)=r_{n}(t;\epsilon)\,\epsilon^{2},$
for some coefficient function, $r_{n}$. Every quantity on the left can be explicitly computed from the series (4) and the tabulated values of $C_{n}$. Therefore the values of $r_{n}(t;\epsilon)$ can be computed for each $\epsilon$. By Taylor analysis, the limiting expressions as $\epsilon\to 0$ are found to be
(15) $r_{n}(t;0)=-2C_{n}E^{\prime}(2^{1/n}t;2^{1/n})t-\big{(}C_{n}E(2^{1/n}t;2^{1/n})\big{)}^{2}.$
These functions are shown for a few values of $n$ in Figure 2. Evidently, the $r_{n}(t;0)$ are bounded for all $t\geq 0$, which provides evidence that the approximation (13) satisfies (7) uniformly to ${\rm O}(\epsilon^{2})$. The implied constant grows as $\alpha\to 1$, however, as can be seen by examining the vertical scales in Figure 2. Therefore, the usefulness of (13) as supplier of initial guesses to the numerical method again diminishes accordingly as $n$ increases.
Figure 2: The coefficient functions $r_{n}(t;\epsilon)$ as defined by (14), in the limit as $\epsilon\to 0$. The observation that these functions are bounded on $[0,\infty)$ provides support for the validity of the formulas (13) as uniform approximations to the solutions of (7).
The details of the numerical method are postponed to the next section, but let us now solve (2) (via (7)) for values of $\alpha$ close to the characteristic values. A few such solutions are shown in Figure 3, where the perturbation and numerical approximations are displayed. Evidently, the perturbation approximations provide good qualitative descriptions of the solutions.
@@
Figure 3: Numerical solutions $v$ of the perturbation equation (7) (solid curves) in comparison to the perturbation approximations (13) (dashed curves). The scaling constants $C_{1}$, $C_{2}$ and $C_{3}$ are given in Table 1. In each frame the value of $\alpha$ is $\alpha_{n}+\epsilon$, with numerical values of these parameters given in the titles and legends.